How much do you need to invest if the intrest rate is 5% and you need a total of $1800 in 4 years?
Answer by LinnW(1048)

Let x = the amount needed
x*(1+0.05)^4 = 1800
x*(1.05)^4 = 1800
x*(1.2155065) = 1800
divide each side by (1.2155065)
x = 1800/(1.2155065)
x = 1480.86
checking,
1480.86*(1.05)^4 = 1799.994 so we are very close